OpenStudy (openstudygirl2):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. b over fifteen minus one equals twenty-nine Step 1: b over fifteen minus one plus one equals twenty-nine plus one Step 2: b over fifteen equals thirty Step 3: b over fifteen divided by fifteen equals thirty divided by fifteen Step 4: b equals two

OpenStudy (openstudygirl2):

@ganeshie8

OpenStudy (openstudygirl2):

@Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

(b/15) -1 =29 eq is this now add 1 on both side

OpenStudy (openstudygirl2):

Okay But which step did you mess up on and why can you help me find that out :D

OpenStudy (er.mohd.amir):

(b/15)=29+1 (b/15)=30

OpenStudy (openstudygirl2):

ok

OpenStudy (openstudygirl2):

Is the step he mess up on Step 3 ?

OpenStudy (er.mohd.amir):

multiply by 15 on both sides (b/15) *15=30*15

OpenStudy (er.mohd.amir):

b=30*15

OpenStudy (er.mohd.amir):

b=450

OpenStudy (openstudygirl2):

ok

OpenStudy (er.mohd.amir):

so in which step he does goes wrong?

OpenStudy (openstudygirl2):

Step 3 ?

OpenStudy (er.mohd.amir):

yes

OpenStudy (openstudygirl2):

Okay! But i dont understand what i would write for my answer?

OpenStudy (openstudygirl2):

Can you tell me what to Write ?

OpenStudy (er.mohd.amir):

he not divide at step 3 but multiply in that step makes ans correct

OpenStudy (openstudygirl2):

So instead of dividing he should of Multiplied ?

OpenStudy (er.mohd.amir):

yes for removing 15 from denominator from b/15

OpenStudy (openstudygirl2):

Okay thx can i ask you another question ?

OpenStudy (er.mohd.amir):

ok

OpenStudy (openstudygirl2):

Okay Question is : Which measures of central tendency are most appropriate for the data set? Why? zero, four, one, three, four, two, one, zero, two, three, and fifteen.

OpenStudy (er.mohd.amir):

i have 10 min only here please quickly

OpenStudy (openstudygirl2):

You can just give me the Answer if that is better XD

OpenStudy (er.mohd.amir):

ok

OpenStudy (openstudygirl2):

This is another pic of the question if you dont understand XD

OpenStudy (er.mohd.amir):

there are mainly 3 central tendency mean ,median ,mode

OpenStudy (openstudygirl2):

ok

OpenStudy (openstudygirl2):

SO what is my answer ?

OpenStudy (er.mohd.amir):

i thick mean is better than other two

OpenStudy (openstudygirl2):

So then what do i write for my answer and why ?

OpenStudy (er.mohd.amir):

mean =sum of all data/ number of data =0+0+1+1+2+2+3+3+4+4+15/11 =35/11

OpenStudy (openstudygirl2):

ok

OpenStudy (er.mohd.amir):

=3.18

OpenStudy (openstudygirl2):

ok

OpenStudy (er.mohd.amir):

which is more correct since median is middle term which is 2 (written in ascending order 2 in the middle)

OpenStudy (openstudygirl2):

Mode or median

OpenStudy (er.mohd.amir):

and mode is most frequent term which is not available

OpenStudy (openstudygirl2):

so is it median ?

OpenStudy (er.mohd.amir):

so ans is mean is more appropriate for given data

OpenStudy (openstudygirl2):

ok thx